### v1

@@ Abstract
Fully tensorial theory of hypercomplex neural networks is given. The key point is to observe that the algebra multiplication can be represented as a rank three tensor. This approach is attractive for neural network libraries that support effective tensorial operations.
@@ 1 Introduction
The core of NN implementations lies in linear algebra usage. In most popular programming libraries (TensorFlow [1], PyTorch [15]), the most popular architecture is feed forward NN, which is based on a stack of layers where the data passes between them unidirectionally. Optimized tensorial operations realize the flow of the data.
@@
There are different algebraic extensions. One of these paths is Algebraic Neural Networks [14], where the additional endomorphism operations on data are performed. The other algebra-geometry direction is the neural networks based on Geometric Algebra/Clifford algebra [4, 17]. Recently, Parametrized Hypercomplex Neural Networks were invented [10] for convolutional layers. They can learn optimal hypercomplex algebra adjusted to data, exploring optimized Kronecker product. However, in some applications keeping hyperalgebra of even more general algebra parameters as (fixed) hyperparameters is needed. In such a way we can optimize algebra structure at the metalevel.
@@
In this paper we discuss implementation in which we change classical real algebra computations into various hypercomplex or even more general algebras computations. This approach presented, e.g., in [3] is not new. However, there is a revival of interest of this direction due to better complexity properties in such areas as image processing [18] or time series analysis [11]. In these contributions the Open Source code [18] for specific four-dimensional hypercomplex algebras was used.
@@
The implementation explained in this article significantly expands the ideas from [18] for arbitrary algebras, including hypercomplex ones. The algorithms described here agree with the NN presented in [18] for specific 4-dimensional hypercomplex algebras. However, implementation of [18] was obtained by constructing an additional multiplication structure from the multiplication table for the hypercomplex algebra, which is treated as additional step in setting up neural network. Our approach permit us to omit this complexity and generalize to arbitrary algebras. This is very important contribution form theoretical treatment of general algebraic approach to hypercomplex neural networks.
The main contribution of this paper is following:
@@
  • summarize basic concepts on tensorial operations in terms of hypercomplex and more general algebras, especially, we noted that the algebra multiplication can be expressed as a third-rank tensor,
@@
  • provide general algorithm for computations within hypercomplex dense layer,
@@
  • provide general algorithms for 1-, 2-, and 3-dimensional hypercomplex convolutional layer computations.
2 Methods
@@
This section provides an overview of the mathematical theory behind the operations used in implementing hypercomplex neural networks. This is a tenet of methods used in this paper. These are classical notions explained in detail in standard references, e.g., [2].
@@ 2.1 Tensors
The primary object that is used in NN implementations is a tensor. It relies on the tensor product described in the following definition
@@
The tensor product of two vector spaces $V_{1}$ and $V_{2}$ over the field $\mathbb{F}$ is the vector space denoted by $V_{1}\otimes V_{2}$ and defined as a quotient space $V_{1}\times V_{2}/L$, where $L$ is a subspace of $V_{1}\times V_{2}$ that is spanned by
$\begin{array}[]{c}(v+w,x)-(v,x)-(w,x),\\ (v,x+y)-(v,w)-(v,x),\\ (\lambda v,x)-\lambda(v,x),\\ (v,\lambda x)-\lambda(v,x),\end{array}$ (1)
where $v,w\in V_{1}$, $x,y\in V_{2}$, $\lambda\in\mathbb{F}$.
By induction, it can be defined for $k$ vector spaces $\{V_{i}\}_{i=1}^{k}$ and denoted by $V_{1}\otimes\ldots\otimes V_{k}$ that can be dented by $\otimes_{i=1}^{k}V_{i}$.
The tensor product can also be defined for duals spaces $V_{i}^{*}$ of a vector spaces $V_{i}$ for $i\in\{1,\ldots k\}$, $0<k<\infty$, and we can define mixed tensor product made for vector spaces and their duals.
The tensor product of vector space is a vecotr space, so we can define a base, e.g., if the base of $V$ is $\{e_{i}\}_{i=1}^{n_{1}}$ and $W$ is $\{f_{i}\}_{i=1}^{n_{2}}$, then the base of $V\otimes W$ is $\{e_{i}\otimes f_{j}\}_{i,j=1}^{n_{1},n_{2}}$.
The tensor space can be used to decompose any multilinear mapping. It is expressed in the universal factorization theorem for tensor product. It states that for a bilinear mapping $F:V\times W\rightarrow X$ of vector spaces $V,W,X$ can be uniquely factorized by a new mapping $\bar{F}:V\otimes W\rightarrow X$ according to Fig. 1.
@@
Figure 1: Unique factorization property of the tensor product.
Then the map $\bar{F}$ is called the tensor. Moreover, in the map $F$, we can move $X$ to the domain of the map, i.e., we can define $\hat{F}:V\times W\times X\rightarrow\mathbb{F}$ instead of $F$, where $\mathbb{F}$ is a field common to the vector spaces.
@@
In the base $\{e_{i}\}$ of $V$ and $\{f_{j}\}$ of $W$ the bilinear mapping $F:V\otimes W\rightarrow\mathbb{R}$ has the form
$F=F^{ij}e_{i}\otimes f_{j},$ (2)
where $F^{ij}\mathbb{R}$ for all $i,j$, are the coefficients of a numerical matrix that is a representation of the multilinear mapping (tensor) $F$ in the fixed base of $V$ and $W$. The matrix collects the components of the tensor in a fixed tensor base.
The matrix $[F_{ij}]$ is implemented in TensorFlow and PyTorch library as a tensor class. The critical difference is that the mathematical tensors have specific properties of transformations under the change of basis of underlying vector spaces. The libraries implement the tensors as a multidimensional matrix of numbers. Moreover, they do not keep the upper (contravariant) or lower (covariant) position of indices.
@@
The linear mapping $A:V\rightarrow W$ can be written as a mapping $\bar{A}\in V^{*}\otimes W\rightarrow\mathbb{R}$ that can be written in the base $\{e^{i}\}$ of $V^{*}$ and $\{f_{i}\}$ of $W$ as [^1] Here we use Einstein summation convention: repeated bottom and top index indicate the summation over the whole range. $A=A_{i}^{j}e^{i}\otimes f_{j}$. The vector space of linear operators is denoted as $L(V,W)$.
For tensors we also often use abstract index notation where we provide only components of the tensor, e.g., $A_{ij}$, understanding them not as fixed base numerical values but as a full tensor $A_{ij}e^{i}\otimes e^{j}$, see [16],
Since the tensor product is a functor in the category of linear spaces [2], therefore, for a linear mapping $A:V\rightarrow W$, we can define the extension of the mapping for a tensor product space $\otimes_{i=1}^{n}A:\otimes_{i=1}^{n}V\rightarrow\otimes_{i=1}^{n}W$ by acting on product as $\otimes A(v_{1}\otimes\ldots\otimes v_{n})=Av_{1}\otimes\ldots\otimes Av_{n}$ and extending by linearity for all combinations.
This is similar behaviour as for the Cartesian product of vector spaces. The Cartesian product is also a functor, and therefore, all the above operations apply in this case.
@@
  • Broadcasting: it is defined on for a linear operator $A:V\rightarrow W$ to be an multilinear extension
@@
    that is $b_{1}^{n}(A)(v_{1},\ldots,v_{n})=(Av_{1},\ldots,Av_{n})$. Similar broadcasting is realized for the tensor product. Both operations relies on functoriality of Cartesian product and tensor product.
  • Transposition/Permutation: The transposition of two components relies on the following fact: there is the unique mapping $\tau:V\otimes W\rightarrow W\otimes V$ that simply reverses the order of factors $\tau(v\otimes w)=w\otimes v$. We can extend it to arbitrary permutation of $n$ numbers, $p:\{1,\ldots,n\}\rightarrow\{1,\ldots,n\}$, we have $\tau_{p}(v_{1}\otimes\ldots\otimes v_{n})=v_{p(1)}\otimes\ldots\otimes v_{p(n)}$. We can define a similar operation for the Cartesian product.
@@
  • Reshaping: For a pair of indices $i,j$ of the range $0\leq R(i)<\infty$ and $0\leq R(j)<\infty$ we can define a reshape operation, which is the new sole index $Rs(i,j)$ that value depends on the values of $i,j$ given by the function: $Rs(i,j)=iR(j)+j$. We can extend the operation for the pair $(p,p+1)$ of neighbour indices of a tensor by $Rs_{p}A_{i_{1}\ldots i_{p}i_{p+1}\ldots i_{k}}=A_{i_{1}\ldots Rs(i_{p},i_{p+1})\ldots i_{k}}$, where $p+1\leq k$. This operation changes only the way of indexing; however, it is useful in applications. In the abstract index notation we can write $Rs_{p}A_{i_{1}\ldots i_{p}i_{p+1}\ldots i_{k}}=A_{i_{1}\ldots i_{p-1}Rs(i_{p},i_{p+1})i_{p+2}\ldots i_{k}}$.
  • Contraction: For a two tensors $A$ and $B$, and a fixed base $\{e_{i}\}$ of $V$, the contraction of indices $p$ (related to $V$) and $q$ (related to $V^{*})$ is (note implicit sum): $C_{p,q}(A,B)=A(\ldots\underbrace{e_{i}}_{p}\ldots)\otimes B(\ldots\underbrace{e^{i}}_{q}\ldots)$ where $A\in\ldots\otimes\underbrace{V}_{p}\otimes\ldots$ and $B\in\ldots\otimes\underbrace{V^{*}}_{q}\otimes\ldots$.
@@
  • Concatenation: joins the tensors $\{A^{(i)}_{i_{1}\ldots i_{l}}\}_{i=1}^{n}$ of the same shape along given dimension $j$, i.e., $K_{j}(\{A^{(i)}\}_{i=1}^{n})=A_{i_{1}\ldots i_{j-1}k_{j}i_{j}\ldots i_{l}}$.
@@
In this part we introduce mathematical concepts related to hypercomplex and general algebras, as in the following definition.
@@
The algebra over a field $\mathbb{F}$ is a vector space $V$ equipped with a product - a binary operation __$\_\cdot\_:V\times V\rightarrow V$ with the following properties:
@@
Moreover, the algebra is commutative if $x\cdot y=y\cdot x$ for $x,y\in V$.
@@
For real numbers $\mathbb{R}$ we have $V=\{e_{0}\}$, $\mathbb{F}=\mathbb{R}$ with $e_{0}\cdot e_{0}=1$.
Complex numbers $\mathbb{C}$ can be obtained from commutative algebra with $V=\{e_{0},e_{1}\}$, $F=\mathbb{R}$ and $e_{0}\cdot e_{0}=e_{0}$, $e_{0}\cdot e_{1}=e_{1}$ and $e_{1}\cdot e_{1}=-e_{0}$.
By convention we will always assume that the neutral element of the algebra mutliplication is $e_{0}$.
Efficient use of algebras in computations based on tensors (TensorFlow, PyTorch) relies on converting the product within the algebra into a tensor operation.
Treating algebra as a vector space with additional structure of vector multiplication [^2] It can be define using a functor that cast category of algebras into the category of linear spaces. we have the following definition, which is essential to the rest of the paper.
Definition 3.
For an algebra $V$ over $F$ the product can be defined as a tensor $A\in V^{*}\otimes V^{*}\otimes V$. Selecting the base of $V=span\{e_{i}\}_{i=0}^{1}$ and the dual base $V^{*}=\{e^{i}\}_{i=1}^{n}$ with $e^{i}(e_{j})=\delta_{j}^{i}$, the product has the form
@@
Then the multiplication table entry is presented in (5).
@@
The tensor coefficients (abstract index notation), $A_{ij}^{k}$, play the same role as structure constants for a group [8]. These coefficients in a fixed base can be represented as a multidimensional matrix (called tensors in TensorFlow and PyTorch). When the algebra is commutative, then $A(x,y)=A(y,x)$ or $A_{ij}^{k}=A_{ji}^{k}$.
3 Results
In this section, we provide mathematical details of the implementation of hypercomplex dense and convolutional neural networks.
We do not distinguish co- and contravariant indices in the description and write them at the bottom level. Moreover, the tensor is treated as a multidimensional array, with indices starting from 0. This is a standard convention in the tensor libraries such as TensorFlow and PyTorch.
We assume that the algebra multiplication structure constant is fixed and stored in the tensor (in abstract index notation) $A=A_{i_{a}j_{a}k_{a}}$, see Subsection 2.2.

### v2

@@ Abstract
Fully tensorial theory of hypercomplex neural networks is given. It allows neural networks to use arithmetic based on arbitrary algebras. The key point is to observe that algebra multiplication can be represented as a rank three tensor and use this tensor in every algebraic operation. This approach is attractive for neural network libraries that support effective tensorial operations. It agrees with previous implementations for four-dimensional algebras.
@@ 1 Introduction
The core of NN implementations lies in linear algebra usage. In most popular programming libraries (TensorFlow [1], PyTorch [17]), the common architecture is the feed-forward NN, which is based on a stack of layers where the data passes between them unidirectionally. Optimized tensorial (multidimensional arrays) operations realize the flow of the data.
@@
There are different algebraic extensions. One of these paths follows Algebraic Neural Networks [16], where the additional endomorphism operations on data are performed. The other algebra-geometry direction is the neural networks based on Geometric Algebra/Clifford algebra [4, 19], starting from the most prominent example of Quaternions [9, 12]. Recently, Parametrized Hypercomplex Neural Networks were invented [11] for convolutional layers. They can learn optimal hypercomplex algebra adjusted to data exploring optimized Kronecker product. However, some applications need hyperalgebra, or even more general algebra parameters, as (fixed) hyperparameters. In such a way we can optimize algebra structure at the metalevel.
@@
In this paper, we discuss implementation in which we change classical real algebra computations into various hypercomplex or even more general algebra computations by including multiplication in the algebra as a tensorial operation. Hypercomplex NN approach is not new, e.g., [3, 9]. However, interest in this direction is revived due to better complexity properties in such areas as image processing [20] or time series analysis [13]. In these contributions the Open Source code [20] for specific four-dimensional hypercomplex algebras were used.
@@
The implementation explained in this article significantly expands the ideas from [20, 9] for arbitrary algebras, including hypercomplex ones. The algorithms described here agree with the NN presented in [20, 9] when limited to 4-dimensional hypercomplex algebras. The 4-dimensional hyperalgebras are useful in encoding RGB color data in images. Moreover, implementation of [20] was obtained by constructing an additional multiplication matrix from the multiplication table for the hypercomplex algebra. This is an additional step in setting up a neural network. Our approach permits us to omit this complexity and generalize to arbitrary algebras since we input only the multiplication table of the algebra as it can be found in standard references, e.g., [2]. Moreover, the limitation to 4-dimensional algebras can be relaxed. This is important for processing data that naturally encodes in $n$-tuples ($n>0$, integer) that can be encoded as a single element of an $n$-dimensional algebra. This is a crucial contribution from the theoretical treatment of the general algebraic approach to hypercomplex neural networks.
The main contribution of this paper is the following:
@@
  • summarize basic concepts of hyperalgebras and more general algebras, and prove that the algebra multiplication can be expressed as a third-rank tensor,
@@
  • provide a general algorithm for computations within the hypercomplex dense layer that generalizes ideas of [20, 9] for arbitrary algebras of arbitrary algebra dimensions,
@@
  • provide general algorithms for 1-, 2-, and 3-dimensional hypercomplex convolutional layer computations that generalizes ideas of [20, 9] for arbitrary algebras of any algebra dimensions.
The paper is organized as follows: after recalling the general algebraic notions in the next section, the algorithms for general architectures of dense and convolutional hypercomplex NN are presented. Finally, in Section 4, we present some conclusions and future directions for work.
2 Theory
@@
This section provides an overview of the mathematical theory behind the operations used in implementing hypercomplex neural networks. This is a tenet of methods used in this paper. A tensor is a classical multilinear algebra notion explained in detail in standard references, e.g., [2]. We rephrase it here to prepare the background for explaining how algebra multiplication can be expressed as a third-rank tensor.
@@ 2.1 Tensors
The primary object used in NN implementations is a ’tensor’. In typical NN libraries such as TensorFlow and PyTorch tensors are synonyms of multidimensional matrix. In algebra, tensors, apart from being a mathematical object that can be represented as multidimensional arrays after selecting a base of a vector space, have many interesting and valuable properties. It relies on the tensor product.
We will consider vector spaces: $V$ and its dual $V^{*}$ over a field $\mathbb{F}$ that can represent, e.g., real numbers. For a fixed base $\{e_{i}\}$ of $V$ the dual space has a base $\{e^{j}\}$, where $e^{i}(e_{j})=\delta^{i}_{j}$, where $\delta^{i}_{j}$ is the Kronecker delta defined as follows $\delta^{i}_{j}=\left\{\begin{array}[]{cc}1&i=j\\ 0&i\neq j\end{array}\right.$.
@@
The tensor product of two vector spaces $V_{1}$ and $V_{2}$ over the field $\mathbb{F}$ is the vector space denoted by $V_{1}\otimes V_{2}$ and defined as a quotient space $V_{1}\times V_{2}/L$, where the subspace $L\subset V_{1}\times V_{2}$ is spanned by
$\begin{array}[]{c}(v+w,x)-(v,x)-(w,x),\\ (v,x+y)-(v,x)-(v,y),\\ (\lambda v,x)-\lambda(v,x),\\ (v,\lambda x)-\lambda(v,x),\end{array}$ (1)
where $v,w\in V_{1}$, $x,y\in V_{2}$, $\lambda\in\mathbb{F}$. The equivalence class of $(v,x)$ in $V_{1}\otimes V_{2}$ is denoted by $v\otimes x$.
From practical point of view, for $v+\lambda w\in V_{1}$ and $x+y\in V_{2}$, $\lambda\in\mathbb{F}$, then $(v+\lambda w)\otimes(x+y)=v\otimes x+v\otimes y+\lambda w\otimes x+\lambda w\otimes y$ belongs to $V_{1}\otimes V_{2}$.
By induction, the tensor product can be defined for $k$ vector spaces $\{V_{i}\}_{i=1}^{k}$ and is denoted by $V_{1}\otimes\ldots\otimes V_{k}$ or shortening notation as $\otimes_{i=1}^{k}V_{i}$.
The tensor product can also be defined for dual spaces $V_{i}^{*}$ of a vector spaces $V_{i}$ for $i\in\{1,\ldots k\}$, $0<k<\infty$, and we can define mixed tensor product made from vector spaces and their duals.
The tensor product of a vector space is a vector space, so we can define a base, e.g., if the base of $V$ is $\{e_{i}\}_{i=1}^{n_{1}}$ and $W$ is $\{f_{i}\}_{i=1}^{n_{2}}$, then the base of $V\otimes W$ is $\{e_{i}\otimes f_{j}\}_{(i,j)=(1,1)}^{(n_{1},n_{2})}$.
The tensor product of vector spaces is useful in decomposing any multilinear mapping. It is expressed in the so-called universal factorization theorem for tensor product [2]. For a bilinear mapping $F:V\times W\rightarrow\mathbb{F}$ of vector spaces $V,W$, it states that $F$ can be uniquely factorized by a new mapping $f:V\otimes W\rightarrow\mathbb{F}$, which is expressed by the commutative diagram in Fig. 1.
@@
Figure 1: Unique factorization property of the tensor product means that $F=f\circ\otimes$.
Then, the map $f$ is called the tensor. This theorem can be generalized to any multilinear mapping.
Definition 2.
A multilinear mapping $F:\underbrace{V\times\ldots\times V}_{p}\times\underbrace{V^{*}\times\ldots\times V^{*}}_{q}\rightarrow\mathbb{F}$ of a vector space $V$ and its dual $V^{*}$, can be factorized by the tensor $f:\otimes_{i=1}^{p}V\otimes\otimes_{j=1}^{q}V^{*}\rightarrow\mathbb{F}$ by $F=f\circ\otimes$.
For fixed bases $\{e_{i}\}$ of $V$ and its dual $\{e^{j}\}$ for $V^{*}$ we can decompose $f$ as [^1] Here we use Einstein summation convention: repeated bottom and top index indicate the summation over the whole range.
$f=f_{i_{1}\ldots i_{p}}^{j_{1}\ldots j_{q}}e^{i_{1}}\otimes\ldots\otimes e^{i_{p}}\otimes e_{j_{1}}\otimes\ldots\otimes e_{j_{q}},$ (2)
where numbers $f_{i_{1}\ldots i_{p}}^{j_{1}\ldots j_{q}}$ are components of the tensor $f$. The tensor $f$ is called $p$-covariant and $q$-contravariant or $(p,q)$ tensor.
The tensor components can be arranged in a multidimensional array and are called ’tensor’ in packages such as TensorFlow or PyTorch. We can also adopt the concept of a tensor’s shape, which is a tuple containing the range of all indices. The nomenclature in these libraries names the axis each index, which represents a vector space in the tensor product.
@@
In the bases $\{e_{i}\}$ of $V$ and $\{f_{j}\}$ of $W$ the bilinear mapping $F:V\times W\rightarrow\mathbb{R}$ factorizes through tensor product space as
$F(e_{i},f_{j})=f(e_{i}\otimes f_{j})=F_{ij},$ (3)
where $F_{ij}\in\mathbb{R}$ for all $i,j$, are the coefficients of a numerical matrix that is a representation of the bilinear mapping (tensor) $F$ in the fixed base of $V$ and $W$. The matrix collects the components of the tensor in a fixed tensor base. We can therefore write tensor in the base $\{e^{i}\}$ of $V^{*}$ and $\{f^{j}\}$ of $W^{*}$, i.e., $e^{i}(e_{j})=\delta^{i}_{j}=f^{i}(f_{j})$, as
$f=F_{ij}e^{i}\otimes f^{j}.$ (4)
We have for two vectors $u=u^{i}e_{i}$, $x=x^{i}f_{i}$
$F(u,x)=f(u\otimes x)=F_{ij}e^{i}\otimes f^{j}(u\otimes x)=F_{ij}e^{i}(u)f^{j}(x)=F_{ij}u^{i}x^{j}.$ (5)
The tensor $f$ has the shape $(dim(V),dim(W))$.
Let us notice that the matrix $[F_{ij}]$ is implemented in TensorFlow and PyTorch library as a tensor class. The critical difference is that mathematical tensors have specific properties of transformations under the change of basis in the underlying vector spaces. The libraries implement the tensors as a multidimensional matrix of numbers. Moreover, the libraries usually do not keep the upper (contravariant) or lower (covariant) position of indices.
@@
The linear operator $F:V\rightarrow W$ can be written as a mapping $f\in V^{*}\otimes W\rightarrow\mathbb{R}$ with partial application of a vector. It can be written in the base $\{e^{i}\}$ of $V^{*}$ and $\{f_{i}\}$ of $W$ as $f=F_{i}^{j}e^{i}\otimes f_{j}$, where $F_{i}^{j}$ is the matrix of $F$ in the mentioned bases of $V$ and $W$. The vector space of linear operators is denoted as $L(V,W)$. One can prove that there is a bijection between linear operators and mixed rank $2$ tensors.
For tensors we also often use abstract index notation where we provide only components of the tensor, e.g., $A_{ij}$, understanding them not as fixed base numerical values but as a full tensor $A_{ij}e^{i}\otimes e^{j}$ (cf. [18]).
Since the tensor product is a functor in the category of linear spaces [2], therefore, for a linear mapping $f:V\rightarrow W$, we can define the extension of the mapping for a tensor product space $\otimes_{i=1}^{n}f:\otimes_{i=1}^{n}V\rightarrow\otimes_{i=1}^{n}W$ by acting on product as
$\otimes f(v_{1}\otimes\ldots\otimes v_{n}):=fv_{1}\otimes\ldots\otimes fv_{n},$ (6)
and extending by linearity for all combinations. This is similar rule as for extending linear operators to the Cartesian product of vector spaces. This results from the fact that the Cartesian product is also a functor, just as a tensor product is.
@@
  • Broadcasting: it is defined for a linear operator $A:V\rightarrow W$ to be a multilinear extension to the Cartesian product of spaces
@@
    that is $b_{1}^{n}(A)(v_{1},\ldots,v_{n})=(Av_{1},\ldots,Av_{n})$. Similar broadcasting is realized for the tensor product. Broadcasting relies on the functionality of the Cartesian product and tensor products.
  • Transposition/Permutation: The transposition of two components relies on the following fact: there is isomorphism $\tau:V\otimes W\rightarrow W\otimes V$ that simply reverses the order of factors $\tau(v\otimes w)=w\otimes v$. We can extend it to arbitrary permutation of $n$ numbers, that is, a bijective mapping: $p:\{1,\ldots,n\}\rightarrow\{1,\ldots,n\}$. We define
    $\tau_{p}(v_{1}\otimes\ldots\otimes v_{n})=v_{p(1)}\otimes\ldots\otimes v_{p(n)}$ (8)
    We can define a similar operation for the Cartesian product.
@@
  • Reshaping: For a pair of indices $i,j$ of the range $0\leq R(i)<\infty$ and $0\leq R(j)<\infty$, where $R$ denotes the maximal value of its argument, which is an index. We can define a reshape operation $Rs$, which is the new index $Rs(i,j)$ that value depends on the values of $i,j$ given by the function
    $Rs(i,j)=iR(j)+j.$ (9)
    We can extend the operation for the pair $(p,p+1)$ of neighbor indices of a tensor, using abstract index notation, by $Rs_{p}A_{i_{1}\ldots i_{p}i_{p+1}\ldots i_{k}}=A_{i_{1}\ldots Rs(i_{p},i_{p+1})\ldots i_{k}}$, where $p+1\leq k$. This operation changes only the way of indexing; however, it is useful in applications.
  • Contraction: For a two tensors $A\in\ldots\otimes\underbrace{V}_{p}\otimes\ldots\rightarrow\mathbb{F}$ and $B\in\ldots\otimes\underbrace{V^{*}}_{q}\otimes\ldots\rightarrow\mathbb{F}$, and a fixed base $\{e_{i}\}$ of $V$ and $\{e^{j}\}$ of $V^{*}$, the contraction of indices $p$ (related to $V$ in $A$) and $q$ (related to $V^{*}$ in $B$) is (note implicit sum):
    $C_{p,q}(A,B)=A(\ldots\underbrace{e_{i}}_{p}\ldots)\otimes B(\ldots\underbrace{e^{i}}_{q}\ldots),$ (10)
    or in abstract index notation
    $C_{p,q}(A,B)=A_{\ldots\underbrace{i}_{p}\ldots}\otimes B^{\ldots\overbrace{i}^{q}\ldots}.$ (11)
@@
  • Concatenation: joins the tensors $\{A^{(i)}_{i_{1}\ldots i_{l}}\}_{i=1}^{n}$ of the same shape along given dimension $j$, i.e.,
    $K_{j}(\{A^{(i)}\}_{i=1}^{n})=A_{i_{1}\ldots i_{j-1}k_{j}i_{j}\ldots i_{l}}=A^{(k_{j})}_{i_{1}\ldots i_{j-1}i_{j}\ldots i_{l}}.$ (12)
@@
In this part we introduce mathematical concepts related to hypercomplex and general algebras. We start from
@@
Algebra over a field $\mathbb{F}$ is a vector space $V$ equipped with a product - a binary operation __$\_\cdot\_:V\times V\rightarrow V$ with the following properties:
@@
The algebra is commutative if $x\cdot y=y\cdot x$ for $x,y\in V$.
We can define multiplication in the algebra by defining it on the base $\{e_{i}\}_{i=1}^{dim(V)}$ of underlying vector space $V=span(\{e_{i}\})$.
@@
For real numbers $\mathbb{R}$ we have $V=span(\{e_{0}\})$, $\mathbb{F}=\mathbb{R}$ with $e_{0}\cdot e_{0}=e_{0}$. We generally associate $e_{0}$ with $1\in\mathbb{R}$.
Example 4.
Complex numbers $\mathbb{C}$ can be obtained from commutative algebra with $V=span(\{e_{0},e_{1}\})$, $F=\mathbb{R}$ and $e_{0}\cdot e_{0}=e_{0}$, $e_{0}\cdot e_{1}=e_{1}$ and $e_{1}\cdot e_{1}=-e_{0}$. We associate $e_{0}$ with $1$, and $e_{1}$ with the imaginary unit $i$.
By convention we will always assume that the neutral element of the algebra multiplication, denoted by $1$, is $e_{0}$.
Efficient use of algebras in computations based on tensors (in TensorFlow and PyTorch) relies on converting the product within the algebra into a tensor operation.
Treating algebra as a vector space with the additional structure of vector multiplication [^2] It can be defined using a functor that casts the category of algebras into the category of linear spaces. we have the following Theorem, which is essential to the rest of the paper.
Theorem 1.
For an algebra $V$ over $F$ the product can be defined as a tensor $A:V^{*}\otimes V^{*}\otimes V\rightarrow\mathbb{F}$. Selecting the base of $V=span(\{e_{i}\}_{i=0}^{n})$, $n=dim(V)<\infty$, and the dual base $V^{*}=span(\{e^{i}\}_{i=1}^{n})$ with $e^{i}(e_{j})=\delta_{j}^{i}$, the product is defined by the tensor
@@
The multiplication is defined by partial application of vectors $x=x^{i}e_{i},y=y^{i}e_{i}\in V$
$x\cdot y:=A(x,y)=A^{k}_{ij}x^{i}y^{j}e_{k}.$ (14)
The multiplication table entry is presented in [15].
@@
Proof.
For the proof, we must check if the multiplication defined by $A$ fulfills the axioms of multiplication $x\cdot y$ for an algebra.
By linearity of tensor $(x+y)\cdot z=A(x+y,z)=A(x,z)+A(y,z)=x\cdot z+y\cdot z$, and $x\cdot(y+z)=A(x,y+z)=A(x,y)+A(x,z)=x\cdot y+x\cdot z$, for $x,y,z\in V$.
Moreover, if $\alpha,\beta\in\mathbb{F}$ and $x,y\in V$ we have by linearity of tensor $(\alpha x)\cdot(\beta y)=A(\alpha x,\beta y)=\alpha\beta A(x,y)=\alpha\beta(x\cdot y)$, as required. ∎
Corollary 1.
When the algebra is commutative, then $A(x,y)=A(y,x)$ or $A_{ij}^{k}=A_{ji}^{k}$.
Proof.
Since $x\cdot y=y\cdot x$, so $A(x,y)=A(y,x)$, which results in symmetry of tensor indices $A_{ij}^{k}=A_{ji}^{k}$, as required. ∎
The tensor coefficients (abstract index notation), $A_{ij}^{k}$, play the same role as structure constants for a group [8]. These coefficients in a fixed base can be represented as a multidimensional matrix (called tensors in TensorFlow and PyTorch).
Example 5.
For complex numbers $\mathbb{C}$, $V$ has two dimensional base $\{e_{0},e_{1}\}$ with multiplication matrix given in [16].
$\begin{array}[]{c|c|c}\cdot&e_{0}&e_{1}\\ \hline\cr e_{0}&e_{0}&e_{1}\\ \hline\cr e_{1}&e_{1}&-e_{0}\\ \end{array}$ (16)
Then the multiplication can be written as the following tensor
$A_{\mathbb{C}}=e^{0}\otimes e^{0}\otimes e_{0}+e^{0}\otimes e^{1}\otimes e_{1}+e^{1}\otimes e^{0}\otimes e_{1}-e^{1}\otimes e^{1}\otimes e_{0}.$ (17)
3 Neural Networks Architectures
In this section, we provide mathematical details of implementing hypercomplex dense and convolutional neural networks.
We do not distinguish co- and contravariant indices in the description, and we write them in subscripts. Moreover, the tensor is treated as a multidimensional array, with indices starting from 0. This is a standard convention in the tensor libraries such as TensorFlow and PyTorch.
We assume that the algebra multiplication structure constant is fixed and stored in the tensor (in abstract index notation) $A=A_{i_{a}j_{a}k_{a}}$ (cf. Subsection 2.2).
We want to stress out that the proposed algorithms are generalizations of algorithms from [20, 9] from $4$-dimensional algebras to arbitrary dimensional algebras using tensor operations that are optimized for tensor operations from packages like TensorFlow or PyTorch. They are implemented in Hypercomplex Keras Python library [22] and described elsewhere [15]. The implementation was constructed to agree for $4$-dimensional algebras with the results from [20].